In a right triangle the hypotenuse is usually labeled c and the legs are a and b.Find the missing side in each of the following.If the answer is not exact, express it to the nearest tenth.
a=7 b=? c=25

Answer by uknouluvme89(30)

a^2 + b^2 = c^2
7^2 + b^2 = 25^2
49 + b^2 = 625
b^2 = 576
b = 24